OpenStudy (anonymous):

A BALL IS DROPPED FROM HEIGHT 40M.DRAW GRAPH FOR A-T,V-T,S-TFOR FIRST TWO BOUNCES?(BALL BOUNCE WITH SAME SPEED AND G=10)

OpenStudy (anonymous):

so what is the question asking you to do, just draw a graph

OpenStudy (anonymous):

i dn't know how to draw the ball falling graph

OpenStudy (anonymous):

are you dealing with Velocity and Acceleration etc..

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

so V=D/T correct? F=M/A os A=F/M correct?

OpenStudy (anonymous):

ok